OpenStudy (anonymous):

Solve for t in d = rt, if d = 57 and r = 30

OpenStudy (anonymous):

t= d/r

OpenStudy (anonymous):

d divided by r?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

huhh

OpenStudy (anonymous):

d = rt d = 57 r = 30 57 = 30(t) t = 57/30 t = 1.9